x=-1/2
f(x)=-1/2x+1

Answer by checkley77(12844)

x=-1/2
f(x)=-1/2x+1
f(-1/2)=-1/2*-1/2+1
f(-1/2)=1/4+1
f(-1/2)=1 1/4 or 5/4 or 1.25 Answer.